A particle has an initial speed of 27 m/s rightward from the origin at t-0. It decelerates at 6t m/s2 (t in seconds). Find the speed after it has traversed 10m of distance. At what time does this occur 1. 2. At t = 0 a particle has a speed of 8 m/s rightward and is located at s-10m. Knowing a 2v3 (v in m/s) find the location and velocity at 4 seconds. What is the maximum arrow speed if its acceleration decreases to zero linearly from A to B. The release value at A is 192,000 in/s2. Determine the time from A to B. 3. Bl 24

Please answer all 3. Thanks
Answer #1

kmot shat dt rothd七1) 12-11 = 丨一3E.dt t ) 22.d t 2ースニ 2.at tazo-372 sec 2) S 3 d七1(3) 856 2-6 七 2-56 ごニ七ナ1 0 2. d) t -t七十 ds d t 七t2. 2 2+11.38623 25万 -32-バ 2- -ts1386 3825 So 932s +I1 38 6 25 S = 12, s 79 sm at am -191000 ㄦ 2dt V ơn beth 4 x d七 七= dx七 89-11 2七 39 41 24 399y t-
Answer #1

kmot shat dt rothd七1) 12-11 = 丨一3E.dt t ) 22.d t 2ースニ 2.at tazo-372 sec 2) S 3 d七1(3) 856 2-6 七 2-56 ごニ七ナ1 0 2. d) t -t七十 ds d t 七t2. 2 2+11.38623 25万 -32-バ 2- -ts1386 3825 So 932s +I1 38 6 25 S = 12, s 79 sm at am -191000 ㄦ 2dt V ơn beth 4 x d七 七= dx七 89-11 2七 39 41 24 399y t-